# Geodetic Exploration of the Interior Structure of Enceladus

### v1

@@ 2.3 Physical libration
Owing to Enceladus’s moderate orbital eccentricity and slightly elongated shape, it is subject to periodic torques that force harmonic oscillations in its orientation, called physical librations. This libration is superimposed on the overall synchronous orbit of the moon around Saturn. The amplitude of physical libration depends mainly on the moments of inertia of the body, and whether the outer surface layer–the ice shell–is coupled to the deep interior (e.g., Rambaux et al., 2010), and to less extent to the rigidity of the shell. The estimated libration of Enceladus based on the analysis of observations made by Cassini is too large to be explained by an ice shell that is mechanically coupled to the core (Thomas et al., 2016). Recent work by Park et al. (2024) suggests a slightly lower amplitude of libration, although still higher than expected if the ice shell and the core are coupled. A global ocean which mechanically decouples the shell from the rocky core is inferred from the libration amplitude (Thomas et al., 2016). It should be noted that the shell and core might decouple without a global ocean in the presence of another frictionless boundary layer, such as found in soft-bedded glacier systems on Earth (Johnston & Montési, 2017). While the scenario of a non-global ocean cannot be ruled out by the constraint from libration, here we assume the presence of a global subsurface ocean.
@@
$\gamma_{I}=\frac{1}{C_{c}C_{s}}\frac{4e[K_{3}K_{5}+K_{2}K_{6}-n^{2}C_{c}K_{3})]}{(n^{2}-\sigma_{1}^{2})(n^{2}-\sigma_{2}^{2})},$ (7)
where $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ describe the torques between the planet, the shell, and the core, $C_{c}$ and $C_{s}$ are the polar moments of inertia of the core and the shell, $\gamma_{c}$ and $\gamma_{I}$ are the amplitudes of libration of the core and the shell, and $\mathcal{M}$ is the mean anomaly, and
@@
Equation (8) represents the natural frequencies of free libration of the body (Van Hoolst et al., 2013). In order to compute the flattening of each of the layers, as required for the calculation of the torques between layers in equations (78), we use the approach of Tricarico (2014) that provides hydrostatic equilibrium figures for moons in eccentric orbits, along with the global shape from the best-fitting ellipsoid model suggested by Park et al. (2024).
@@ 2.4 Static gravity field
Current gravity data for Enceladus is limited to that obtained from three Cassini fly-bys. Therefore, we are limited to gravity field coefficients of spherical harmonic degree $l=2,3$. To further constrain the interior structure, we use the moment of inertia and the $l=3$ zonal potential coefficients ($J_{3}$). We do not use the $l=2$ gravity coefficients ($J_{2}$ and $C_{22}$) because of the underlying assumptions required to account for the non-hydrostatic component of the gravity field and the compensation mechanism, as discussed in Hemingway & Mittal (2019). Direct measurement of the moment of inertia of Enceladus is possible only by measuring the precession of the body. Hemingway et al. (2018) propose to derive the moment of inertia from long-wavelength $l=2$ gravity-topography admittance measurements ($Z_{22}$ and $Z_{20}$). We refer the reader to McKinnon (2015), Hemingway et al. (2018), and Hemingway & Mittal (2019) for the details of this derivation. Here, we only consider the moment of inertia when we explore the potential impact of future observations. To compute the zonal non-hydrostatic degree-3 static gravity coefficient ($J_{3}$), we follow an approach similar to that described in Genova et al. (2024) to compute the zonal harmonic gravity coefficient. The gravity-topography admittance is computed as (Turcotte et al., 1981):
@@
$H_{30}^{nhyd}$ is the non-hydrostatic topography coefficient. This method implies an unconstrained degree of compensation. Following Genova et al. (2024), we assume Airy compensation, implying $C_{l}^{0}=1$. This assumption does not affect the uncertainties we derive for the observations and model parameters.
@@ 3 End-member scenarios for tidal dissipation in Enceladus
The viscosity of consolidated rock is higher than $10^{18}$ Pa.s (e.g., Hirth & Kohlstedt, 2003). However, unconsolidated material, or even a mixture of rock and ice, may have a significantly lower effective viscosity. In this spirit, Choblet et al. (2017) propose a range of lower core viscosities that allows substantial tidal heating in the core of Enceladus. Alternatively, tidal dissipation in the icy crust may explain the tidal heating at Enceladus if the temperature of ice is sufficiently warm, and thus the effective shell viscosity is sufficiently low. The estimated tidal heating from these different scenarios is shown in Figure 1(a). The four narrow regions determined by the red lines indicate the ranges of the estimated tidal heating in Enceladus and the associated effective viscosities of the shell and the core, assuming that the tidal heating is fully balanced by the heat loss rate at the surface in the range of 15–40 GW (Howett et al., 2011; Hemingway & Mittal, 2019; Nimmo et al., 2023; Park et al., 2024). We show the variation in tidal heating with the viscosity of the shell (core) in Figure 1 panels (b) and (c) for a fixed value of core (shell) viscosity. Dissipation peaks at certain viscosities of both the shell and core at which the associated Maxwell relaxation time is close to the period of excitations, and the tidal response transitions from anelasticity-dominated behavior to viscous-dominated behavior. Therefore, the estimated tidal dissipation rate is compatible with two regions around the peak dissipation (highlighted in gray). For the interior structure assumed here, we find that an effective core viscosity O(1013) Pa.s can satisfy the total estimated heat, consistent with that suggested by Choblet et al. (2017).
@@
The two end-member scenarios manifest themselves differently in the absolute values of $k_{2}$, $h_{2}$, and $l_{2}$ and their respective phase lags (Figure 2). When the effective viscosity of the rocky core, $\eta_{core}$, is $\lesssim 10^{13}$ Pa.s., deformation of the core will significantly affect the amplitude of  $k_{2}$ and its phase lag $\epsilon_{k_{2}}$ (Figure 2 panels (a) and (d)). However, tidal deformation of the core does not substantially affect tidal displacement at the surface, resulting in only a small change in $h_{2}$. This small influence on $h_{2}$ is found to be less than 10% for Enceladus. The compatibility of the steady-state heat budget with both dissipation scenarios presents a degeneracy that requires additional constraints to resolve. In the case of Enceladus and for most of the icy bodies where the ice shell is mechanically decoupled from the solid interior by an ocean, the ambiguity in the location of the tidal heating can be resolved by considering the tidal response at the surface and the associated phase lag. In the case of a mechanically coupled interior or tidally inactive interior layers, there is a strong correlation between the amplitude and phase lag of deformations ($h_{2}$, $l_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{l_{2}}$) on the surface and those of the observed gravitational signal. However, a lack of such a correlation would imply that tidal activity in deeper interior layers decoupled from the surface, such as the core, is causing a perturbation in the gravitational field of the body. In other words, if the ice shell is the primary locus of tidal heat generation, then the shell undergoes considerably larger deformations with a significant phase lag that incorporates viscoelastic deformations, as opposed to the case where the tidal activity occurs predominantly in the rocky core and the deformations in the shell remain largely elastic. For a given interior structure model of Enceladus, the tidal deformations on the surface in the two cases differ by approximately a factor of two (Figure 2 panels (b) and (c)). The phase lags associated with tidal displacements can differ by tens of degrees between the two cases.
@@
Ermakov et al. (2021) and Genova et al. (2024) propose suites of geophysical measurements to study the interior of Enceladus. However, the ranges of elastic and viscoelastic parameters they consider for the rocky core is such that tidal deformation and dissipation in the core are effectively negligible. Ermakov et al. (2021) assume $\eta_{core}>10^{20}$ Pa.s, $\mu_{core}>10$ GPa , and Genova et al. (2024) assume $\eta_{core}>10^{16}$ Pa.s, $\mu_{core}>50$ GPa. The results of these studies represent only one of the tidal dissipation scenarios, i.e., tidal dissipation occuring in the shell without effective contribution from the core. Here, we broaden the parameter space to allow for the possibility of significant tidal dissipation in the core. Note that the commonly assumed viscosity of ice, as suggested by Goldsby & Kohlstedt (2001) ($\eta_{\rm ice}\approx 10^{14}$ Pa.s,) corresponds to ice somewhat near its melting point. Due to the strong temperature dependence of viscosity and very low temperature conditions at Enceladus that may persist deep within the shell, assuming a uniform viscosity of $10^{14}$ Pa.s for the shell (Efroimsky, 2018; Ermakov et al., 2021; Genova et al., 2024) is unrealistic.
@@
Figure 1: (a) Tidal heating (in GW) versus the effective viscosity of the ice shell and the core. The narrow regions bounded by the red lines indicate the range of heat loss that covers the estimates by Howett et al. (2011); Hemingway & Mittal (2019); Park et al. (2024). (b) Tidal heating variation with ice shell effective viscosity for a fixed core viscosity of 1020 Pa.s. (c) Tidal heating variation with core effective viscosity for a fixed shell viscosity of 1016 Pa.s. The gray areas highlight the region where both viscosity and the tidal heating values are compatible with the observed steady state heating. The results are presented for $D_{\rm shell}=14~{}$km, $D_{\rm ocean}=34$ km, $\rho_{\rm ocean}=1000~{}\rm kg/m^{3}$, $R_{\rm core}=204$ km, $\rho_{\rm shell}=930~{}\rm~{}kg/m^{3}$, $\mu_{\rm shell}=3$ GPa, $\kappa_{\rm shell}=10$ GPa, $\rho_{\rm core}=2175~{}\rm~{}kg/m^{3}$, $\kappa_{\rm core}=20$ GPa, $\mu_{\rm core}=15$ GPa. We have assumed an Andrade viscoelastic model.
@@
As described in Section 2.1, there exist various viscoelastic models based on theoretical and laboratory results. Since the nominal values of the effective relaxation time of ice could be close to the orbital period (tidal forcing period) of Enceladus, we examine the choice of Maxwell and Andrade viscoelastic models by computing the tidal response components (Love numbers and heat) using both models. Our analysis shows that if the viscosity of the ice shell is sufficiently high due to the low temperature ($\eta_{\rm shell}\gtrsim 10^{18}$ Pa.s), its relaxation time will be significantly longer than the forcing period, implying a predominantly elastic behavior of the shell, and therefore the choice of the viscoelastic model does not significantly affect the response (Figure 16 in Appendix A). However, if the viscosity of the ice shell is such that the relaxation time ($\tau_{M}~{}=~{}\eta_{\rm shell}/\mu_{\rm shell}$) is close to the forcing period, that is, $\eta_{\rm shell}~{}\approx~{}10^{13-14}$ Pa.s, the Maxwell and Andrade viscoelastic models predict very different surface deformation ($\sim$3 m, see Figure 16 in Appendix A). The two viscoelastic models differ more significantly in their predicted imaginary component of the response, and thus the generated frictional heat, as opposed to the real part of the response which is largely determined by the rigidity of the body. The difference in total tidal heating generated in both the ice shell and the core, following the difference in the $k_{2}$ amplitude and the phase, can reach several GW. Although we do not rule out the applicability of the Maxwell viscoelastic model (see e.g., Hussmann et al., 2016), for further calculations in this paper, we adopt the Andrade model because of its capability to consider transitional anelastic mechanisms missing in the Maxwell model. The extent to which Enceladus’s tidal response deviates from a Maxwell body to an Andrade or more complex viscoelastic body is determined by the large-scale rheological state of its interior (see e.g., Bagheri et al. (2022b); Bierson (2024) for discussions on various viscoelastic models). We investigate the sensitivity of Enceladus’s tidal response to the choice of Andrade parameters in Appendix A.
4 Sensitivity of the ice shell thickness to physical libration and tidal coupling
Using the formulation provided in Section 2.3 and the same physical parameters used in Figures 1 and 2, we consider the sensitivity of the amplitude of libration to the thickness of the ice shell and the effective viscosity of the shell and the core (Figure 3). We use a range for the amplitude of the libration of 360 m$<\gamma_{I}<$750 m, which spans the estimates of Nadezhdina et al. (2016), Thomas et al. (2016), and Park et al. (2024). The sensitivity of the libration to the shell thickness is not uniform for different ice shell thickness ranges. Much higher variations in the libration amplitude are observed when the mean thickness of the ice shell is close to 5–10 km. At these thicknesses, the natural frequency of Enceladus approaches the frequency of the physical libration, resulting in a resonance that significantly enhances the libration amplitude. All current estimates of the physical libration of Enceladus are lower than those suggesting nearly resonant behavior, which rules out a mean ice shell of thickness 5–10 km. The libration amplitude is less sensitive to the viscosity of the ice shell compared to its thickness. However, for ice shell viscosities near the nominal relaxation viscosity of the ice, the effect of the shell’s viscoelastic deformation should not be ignored. This effect is approximately 10% if the viscosity of the shell is as low as $\sim$12.5$\times$1013 Pa.s compared to high-viscosity ice ($\gtrsim 10^{15}$ Pa.s) where there is negligible impact of anelasticity on the predicted libration. As expected, libration is much less sensitive to the viscosity of the core because the only effect from the core is through the gravitational perturbations caused by the core’s tidal deformation on the ice shell, which are insignificant even in the case of a tidally active core.
@@ 4 Sensitivity of the ice shell thickness to physical libration and tidal coupling
Figure 3: Variation of the amplitude of the libration as a function of ice shell thickness and the effective viscosities of shell (a) and the core (b). In panel (a), the core viscosity is assumed equal to $\rm 10^{20}~{}Pa.s$. The region between the red lines shows the libration amplitude spanning the estimates by Thomas et al. (2016); Nadezhdina et al. (2016); Park et al. (2024), i.e., ($\rm 360~{}m<\gamma_{I}<750~{}m)$. Other interior structure parameters are the same as in Figure 2. At $\rm D_{~{}shell}\approx 5$ km, the natural free libration period of the shell approaches the orbital period of Enceladus. The x-axes and the libration color scale are logarithmic.
@@
Radius $R_{E}$ km     251.99${}^{(1)}$
Mass (×$\times 10^{11}$) $M$ kg $1.08101\pm 1.08~{}^{(2)}$
Flattening coeff.(×$\times 10^{6}$) $C_{20}$   –5564.4 $\pm$ 37.58${}^{(1)}$
$-5526.1\pm 35.5~{}^{(3)}$
Quadruple coeff.(×$\times 10^{6}$) $C_{22}$     1542.0 $\pm$ ${}^{(1)}$
@@
Degree-3 coeff.(×$\times 10^{6}$) $J_{3}$ $-180.64\pm 33.95~{}^{(1)}$
  –118.2 ±$\pm 23.5~{}^{(3)}$
@@
Physical libration $\gamma_{I}$ deg         $0.091\pm 0.009~{}^{(1)}$
@@
Orbital mean motion (×$\times 10^{-5}$) $\omega$ rad/s        5.307${}^{(10)}$
Table 1: Existing measurements for Enceladus. The subscripts indicate the references for the quoted values: 1Park et al. (2024), 2Jacobson (2022), 3Hemingway et al. (2018), 4Howett et al. (2011), 5Nimmo et al. (2023), 6Hemingway & Mittal (2019), 7Thomas et al. (2016), 8Nadezhdina et al. (2016), 9Nimmo et al. (2018), 10Roberts & Nimmo (2008). Static gravity coefficients have been adjusted to a reference radius of 251.99 km with respect to the calculated values in Park et al. (2024) that used a radius of 256.6 km.
@@
Shell thickness $D_{\rm shell}$ km $2-50$
@@
Shell shear modulus $\mu_{\rm shell}$ GPa $2-4$
@@
Ocean density $\rho_{\rm ocean}$ kg/m3 $930-1400$
Ocean thickness $D_{\rm ocean}$ km $1-130$
@@
Core radius $R_{\rm core}$ km $120-220$
Core shear modulus $\mu_{\rm core}$ GPa $1-50$
@@
Core density $\rho_{\rm core}$ kg/m3 $2000-2800$

### v3

@@ 2.3 Physical Libration
Owing to Enceladus’s moderate orbital eccentricity and slightly elongated shape, it is subject to periodic torques that force harmonic oscillations in its orientation, called physical librations. This libration is superimposed on the overall synchronous orbit of the moon around Saturn. The amplitude of physical libration depends mainly on the moments of inertia of the body, and whether the outer surface layer, i.e., the ice shell, is coupled to the deep interior (e.g., Thomas et al., 2016), and to a lesser extent to the rigidity of the shell. The estimated libration of Enceladus based on the analysis of Cassini observations is too large to be explained by an ice shell that is mechanically coupled to the core, implying a global ocean (Thomas et al., 2016). Recent work by Park et al. (2024) suggests a slightly lower amplitude of libration, although still higher than expected if the ice shell and the core are coupled. We note that the shell and core might decouple without a global ocean in the presence of a frictionless interface, such as found in soft-bedded glacier systems on Earth (Johnston & Montési, 2017). While the scenario with no global ocean cannot be ruled out by the constraint from libration, here we assume such an ocean is present.
@@
$\gamma=\frac{1}{C_{c}C_{s}}\frac{4e[K_{3}K_{5}+K_{2}K_{6}-n^{2}C_{c}K_{3}]}{(n^{2}-\sigma_{1}^{2})(n^{2}-\sigma_{2}^{2})},$ (7)
where $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ describe the torques between the planet, the shell, and the core, $C_{c}$ and $C_{s}$ are the polar moments of inertia of the core and the shell, $\gamma$ is the amplitudes of libration of the shell, $\mathcal{M}$ is the mean anomaly, and
@@
Equation (8) represents the natural frequencies of free libration of the body (Van Hoolst et al., 2013). To compute the flattening of each of the layers, as required for the calculation of the torques between layers in equations (78), we use the approach of Tricarico (2014) that provides hydrostatic equilibrium figures for moons in eccentric orbits, along with the global shape from the best-fitting ellipsoid model suggested by Park et al. (2024).
@@ 2.4 Static Gravity Field
Current gravity data for Enceladus is limited to that obtained from three Cassini fly-bys. Therefore, we are limited to gravity field coefficients of spherical harmonic degree $l=2,3$. To further constrain the interior structure, we use the moment of inertia and the $l=3$ zonal potential coefficients ($J_{3}$). The moment of inertia is derived from the gravity coefficients ($J_{2}$ and $C_{22}$) with assumptions to account for the non-hydrostatic component of the gravity field and the compensation mechanism, as discussed in Hemingway & Mittal (2019). Directly measuring the moment of inertia of Enceladus is possible only by measuring the precession of the body. Hemingway et al. (2018) propose to derive the moment of inertia from long-wavelength $l=2$ gravity-topography admittance measurements ($Z_{22}$ and $Z_{20}$). See McKinnon (2015), Hemingway et al. (2018), and Hemingway & Mittal (2019) for details. Here, we only consider the moment of inertia when we explore the potential impact of future observations. To compute the zonal non-hydrostatic degree-3 static gravity coefficient ($J_{3}$), we follow an approach similar to that described in Genova et al. (2024). The gravity-topography admittance is computed as (Turcotte et al., 1981; Hemingway & Matsuyama, 2017):
@@
$H_{30}^{nhyd}$ is the non-hydrostatic topography coefficient. We compute $H_{30}^{nhyd}$ by using the formulation of Tricarico (2014) to calculate the hydrostatic shape coefficients, and subtracting them from the observed shape coefficients of Park et al. (2024). This formulation implies an unconstrained degree of compensation. Following Genova et al. (2024), we assume Airy compensation, implying $C_{l}^{0}=1$. This assumption does not affect the uncertainties we derive for the observations and model parameters.
@@ 3 End-member scenarios for tidal dissipation in Enceladus
The viscosity of consolidated rock is higher than $10^{18}$ Pa.s Hirth & Kohlstedt (e.g., 2003). However, unconsolidated material, or even a mixture of rock and ice, may have a significantly lower effective viscosity. In this context, Choblet et al. (2017) propose a range of lower core viscosities that allows substantial tidal heating in the core of Enceladus. Alternatively, tidal dissipation in the icy crust may explain the tidal heating at Enceladus if the temperature of the ice is sufficiently warm and the effective shell viscosity sufficiently low. The estimated tidal heating from these different scenarios is shown in Figure 1(a). The four narrow regions determined by the red lines indicate the ranges of the estimated tidal heating in Enceladus and the associated effective viscosities of the shell and the core, assuming that the tidal heating is fully balanced by the rate of heat loss at the surface in the range of 15–40 GW (Howett et al., 2011; Hemingway & Mittal, 2019; Nimmo et al., 2023; Park et al., 2024). We show the variation in tidal heating with the viscosity of the shell (core) in Figure 1 panels (b) and (c) for a fixed value of core (shell) viscosity. The dissipation peaks at certain viscosities of both the shell and core, at which the associated Maxwell relaxation time is close to the period of excitations and the tidal response transitions from anelasticity-dominated behavior to viscous-dominated behavior. Therefore, the estimated tidal dissipation rate is compatible with two regions around the peak dissipation (highlighted in gray). For the interior structure assumed here, we find an effective core viscosity O(1013) Pa.s can satisfy the total estimated heat, consistent with that suggested by Choblet et al. (2017).
The two end-member scenarios manifest themselves differently in the amplitudes of $k_{2}$, $h_{2}$, and $l_{2}$ and their respective phase lags (Figure 2). When the effective viscosity of the rocky core is $10^{10}~\rm Pa.s\lesssim\eta_{core}\lesssim~10^{13}$ Pa.s., deformation of the core will significantly affect the amplitude of  $k_{2}$ and its phase lag $\epsilon_{k_{2}}$ (Figure 2, panels (a) and (d)). However, tidal deformation of the core does not substantially affect tidal displacement at the surface, resulting in only a small change in $h_{2}$. i.e., less than 10% for Enceladus. Tidal deformations in the core can have a substantial effect on surface deformations ($h_{2}$) if the viscosity of the shell is very low. This scenario is unlikely due to the low temperatures in most of the ice shell.
@@
The compatibility of the produced tidal heat with different regimes of tidal activity in the shell or in the core (Figures 1 and 2) presents an ambiguity in determining the distribution of tidal dissipation that requires additional constraints to resolve. In the case of Enceladus, and for most of the icy bodies where the ice shell is mechanically decoupled from the solid interior by an ocean, the ambiguity in the location of the tidal heating can be resolved by considering the tidal response at the surface and the associated phase lag. In the case of a mechanically coupled interior or tidally inactive interior layers, there is a strong correlation between the amplitude and phase lag of deformations ($|h_{2}|$, $|l_{2}|$, $\epsilon_{k_{2}}$, $\epsilon_{l_{2}}$) on the surface and those of the observed gravitational signal. However, a lack of such a correlation would imply that tidal activity in deeper interior layers that are decoupled from the surface—such as the core—is causing a perturbation in the body’s gravitational field. If the ice shell is the primary locus of tidal heat generation, then it undergoes considerably larger deformations that incorporates viscoelastic deformations and a significant phase lag, as opposed to the case where the deformations in the shell remain largely elastic. For a given interior structure model of Enceladus, the tidal deformations on the surface in the two cases differ by approximately a factor of two (Figure 2 panels b and c). The phase lags associated with the tidal displacements can differ by tens of degrees between the two cases. Section 6 explores the details of determining the primary location of tidal dissipation.
@@
Ermakov et al. (2021) and Genova et al. (2024) propose suites of geophysical measurements to study the interior of Enceladus. However, the ranges of elastic and viscoelastic parameters that they consider for the rocky core are such that the tidal deformation and dissipation in the core are effectively negligible. Ermakov et al. (2021) assume $\eta_{core}>10^{20}$ Pa.s, $\mu_{core}>10$ GPa, and Genova et al. (2024) assume $\eta_{core}>10^{16}$ Pa.s, $\mu_{core}>50$ GPa. The results of these studies implicitly represent only one of the tidal dissipation scenarios, i.e., tidal dissipation occurring in the shell without an effective contribution from the core. Here, we broaden the parameter space to allow for the possibility of significant tidal dissipation in the core. Note that the commonly assumed viscosity of ice, as suggested by Goldsby & Kohlstedt (2001) ($\eta_{\rm ice}\approx 10^{14}$ Pa.s,) corresponds to ice near its melting point. Due to the strong temperature dependence of viscosity and very low temperature conditions at Enceladus that may persist deep within the shell, a uniform viscosity of $10^{14}$ Pa.s for the shell (Efroimsky, 2018; Ermakov et al., 2021; Genova et al., 2024) is unrealistic.
@@
Figure 1: (a) Tidal heating (in GW) versus the effective viscosity of the ice shell and the core. The narrow regions bounded by the red lines indicate the range of heat loss that covers the estimates by Howett et al. (2011); Hemingway & Mittal (2019); Park et al. (2024). (b) Tidal heating variation with ice shell effective viscosity for a fixed core viscosity of 1020 Pa.s. (c) Tidal heating variation with core effective viscosity for a fixed shell viscosity of 1016 Pa.s. The gray areas highlight the region where tidal heating values are compatible with the estimated steady state heating. The results are presented for $D_{\rm shell}=14$~km, $D_{\rm ocean}=34$ km, $\rho_{\rm ocean}=1000~\rm kg/m^{3}$, $R_{\rm core}=204$ km, $\rho_{\rm shell}=930~\rm~kg/m^{3}$, $\mu_{\rm shell}=3$ GPa, $\kappa_{\rm shell}=10$ GPa, $\rho_{\rm core}=2175~\rm~kg/m^{3}$, $\kappa_{\rm core}=20$ GPa, $\mu_{\rm core}=15$ GPa. We assume an Andrade viscoelastic rheology.
@@
As described in Section 2.1, various viscoelastic models exist based on theoretical and laboratory results. Because the nominal values of the effective relaxation time of ice could be close to the orbital period (tidal forcing period) of Enceladus, we examine the choice of Maxwell and Andrade viscoelastic models by computing the tidal response components (Love numbers and heat) using both models. Our analysis shows that if the viscosity of the ice shell is sufficiently high due to the low temperature ($\eta_{\rm shell}\gtrsim 10^{18}$ Pa.s), its relaxation time will be significantly longer than the forcing period, implying a predominantly elastic behavior of the shell, and therefore the choice of the viscoelastic model does not significantly affect the response (See Appendix A). However, if the viscosity of the ice shell is such that the relaxation time ($\tau_{M}~=~\eta_{\rm shell}/\mu_{\rm shell}$) is close to the forcing period, that is, $\eta_{\rm shell}~\approx~10^{13-14}$ Pa.s, the Maxwell and Andrade viscoelastic models predict very different surface deformation ($\sim$3 m, see Appendix A). The two viscoelastic models differ more significantly in their predicted imaginary component of the response, and thus in the generated frictional heat. The real part of the response is largely determined by the rigidity of the body. The difference in total tidal heating generated in the ice shell and the core can reach several GW. Although we do not rule out the applicability of the Maxwell viscoelastic model (see e.g., Hussmann et al., 2016), for further calculations in this paper, we adopt the Andrade model because of its ability to capture transitional anelastic mechanisms missing in the Maxwell model. The extent to which Enceladus’s tidal response deviates from a Maxwell body to an Andrade or more complex viscoelastic body is determined by the large-scale rheological state of its interior (see, e.g., Bagheri et al. (2022b); Bierson (2024) for discussions on various viscoelastic models). We investigate the sensitivity of Enceladus’s tidal response to the choice of Andrade parameters in Appendix A.
4 Sensitivity of physical libration to the shell properties
[…]
@@ 4 Sensitivity of physical libration to the shell properties
Figure 3: Variations of (a) Period of natural (free) libration and (b) amplitude of forced libration of Enceladus as a function of shell thickness and density. The orbital period of Enceladus is shown by gray line in (a). The end-member values of amplitude of libration inferred by Thomas et al. (2016); Nadezhdina et al. (2016); Park et al. (2024) (370 $\gamma<~750 m$) are shown by red lines in (b). Other interior structure parameters are the same as in Figure 2. The x and y axes and the libration amplitude color scale are logarithmic.
[…]
@@
[…]
Mass ($\times 10^{20}$) $M$ kg $1.08101\pm 1.08~^{(2)}$
[…]
$5526.1\pm 35.5~^{(3)}$
[…]
@@
Degree-3 coeff.($\times 10^{6}$) $J_{3}$ $180.64\pm 33.95~^{(1)}$
  –118.2 $\pm 23.5~^{(3)}$
@@
Physical libration $\gamma$ deg     $0.091\pm 0.009~^{(1)}$
@@
[…]
Table 1: Existing measurements for Enceladus. The subscripts indicate the references for the quoted values: 1Park et al. (2024), 2Jacobson (2022), 3Hemingway et al. (2018), 4Howett et al. (2011), 5Nimmo et al. (2023), 6Hemingway & Mittal (2019), 7Thomas et al. (2016), 8Nadezhdina et al. (2016), 9Iess et al. (2014), 10Nimmo et al. (2018), 11Roberts & Nimmo (2008). Static gravity coefficients have been adjusted to a reference radius of 251.99 km with respect to the calculated values in Park et al. (2024) that used a radius of 256.6 km. Heat loss rates in (1), (5), and (6), are estimated from the shell thickness.
@@
Shell thickness $D_{\rm shell}$ km 2502-50
@@
Shell shear modulus $\mu_{\rm shell}$ GPa 242-4
@@
Ocean density $\rho_{\rm ocean}$ kg/m3 9301400930-1400
Ocean thickness $D_{\rm ocean}$ km 11301-130
@@
Core radius $R_{\rm core}$ km 120220120-220
Core shear modulus $\mu_{\rm core}$ GPa 1501-50
@@
Core density $\rho_{\rm core}$ kg/m3 200028002000-2800